A sample of eleven plants gave the following shoot lengths.A simple random sample of 14 people from a certain population gives body mass indices as shown in the table.can we conclude that the BMI is not 35?let significance level be 0.05




1
Expert's answer
2021-11-30T17:02:26-0500

Subject BMI Subject BMI

1 23 8 24

2 25 9 32

3 21 10 57

4 37 11 23

5 39 12 26

6 21 13 31

7 23 14 45


for BMI:

"\overline{x}=30.5,s=10.64"


"H_0:\mu=35" , BMI is 35

"H_a:\mu\neq35" , BMI is not 35


"t=\frac{\overline{x}-\mu}{s/\sqrt n}=\frac{30.5-35}{10.64/\sqrt{14}}=-1.582"


"df=n-1=13"

critical value:

"t_{crit}=2.160"


Since "|t|<t_{crit}" we accept null hypothesis. BMI is 35